3. A software development project at day 70 exhibits an actual cost of $78,000 and a scheduled cost of $84,000 The software manager estimates a value completed of $81,000. What are the cost and schedule variances and CSI? Estimate the time variance.
Answer #1

Given:

AC = $78,000, PV = $84,000, EV = $81,000, AT = 70 days

Cost variance = EV – AC = $81,000 - $78,000 = $3,000

Schedule variance: EV – PV = $81,000 - $84,000 = -$3000

CPI = EV/AC = 1.03

SPI = EV/PV = 0.96

Cost Schedule Index = EV^2/(AC)(PV) = ($81,000)^2/($78,000)($84,000)Cost Schedule Index = 6,561,000,000 / 6,552,000,000

Cost Schedule Index = 1.001

Time Variance = ST –AT = (AT)(CSI)–AT = (70)(1.001)–70

Time Variance = 0.07 days